OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

To determine the slope of a line, we need to either apply the slope formula or draw this out. The slope formula is as follows\[\frac{ (y _{2}-y _{1}) }{ (x _{_{2}}-x_{_{1}}) }\]Here, we simply plug in the coordinates of the two points into their corresponding locations and then solve\[\frac{ 1-(-4) }{ -2-3 }=\frac{ 5 }{ -5 }=-1\]

OpenStudy (anonymous):

im not sure if I'm doing it correctly since i got 5/-5 for my final answer... @graham69 @jdoe0001

OpenStudy (anonymous):

you did it correctly, look at my work

OpenStudy (anonymous):

@chariety101

OpenStudy (anonymous):

oh i see it now thank you so much @graham69

OpenStudy (anonymous):

You're welcome @chariety101